# GPT as Psychologist? Preliminary Evaluations for GPT-4V on Visual Affective Computing

### 09 Mar 2024 v1

@@ Abstract
Multimodal language models (MLMs) are designed to process and integrate information from multiple sources, such as text, speech, images, and videos. Despite its success in language understanding, it is critical to evaluate the performance of downstream tasks for better human-centric applications. This paper assesses the application of MLMs with 5 crucial abilities for affective computing, spanning from visual affective tasks and reasoning tasks. The results show that GPT4 has high accuracy in facial action unit recognition and micro-expression detection while its general facial expression recognition performance is not accurate. We also highlight the challenges of achieving fine-grained micro-expression recognition and the potential for further study and demonstrate the versatility and potential of GPT4 for handling advanced tasks in emotion recognition and related fields by integrating with task-related agents for more complex tasks, such as heart rate estimation through signal processing. In conclusion, this paper provides valuable insights into the potential applications and challenges of MLMs in human-centric computing. The interesting samples are available at https://github.com/LuPaoPao/GPT4Affectivity.
@@ 1 Introduction
The development of multimodal language models (MLMs) has been a topic of growing interest in recent years [22, 12, 23, 43, 21]. MLMs are designed to process and integrate information from multiple modalities, such as text, speech, images, and videos. The development of these models has been driven by the need to improve the accuracy and efficiency of various tasks, such as affective computing, sentiment analysis, and natural language understanding.
MLMs have shown great promise in improving the accuracy and robustness of affective computing systems [37, 27]. These models can process and integrate information from multiple modalities, such as facial expressions, speech patterns, and physiological signals, to infer emotional states accurately [12, 23, 43, 21] with significant implications for various applications, such as healthcare, education, and human-computer interaction.
Figure 1: The propaganda image was generated by Dell2.
Despite the rapid development of MLMs, the need for standardized evaluation metrics is highlighted for accurate assessment. Different from the general language understanding evaluation benchmark that has been widely used to evaluate the performance of language models in NLP tasks [41, 40, 39], similar benchmarks to evaluate the performance of MLMs for affective computing tasks are lacked, which is of great benefit to advance this field.
GPT4-V is a state-of-the-art MLM that has shown remarkable success in various natural language processing tasks [1]. Its ability to process and integrate information from multiple modalities makes it an ideal candidate for evaluating the performance of MLMs in tasks for affective computing. Furthermore, it can invoke a variety of tools that benefit affective computing tasks, such as related program generation with self-correction. For example, GPT4-V can call Dell to generate high-quality visual affective images shown in Fig. 1.
In this paper, we evaluate GPT4-V with 5 typical human-centric tasks, spanning from visual affective tasks and reasoning tasks. We summarize our findings as follows:
(1) GPT4-V is highly accurate in recognizing facial action units. This accuracy can be attributed to its advanced understanding of facial movements and their corresponding emotions, which allows it to effectively identify and analyze facial action units.
(2) GPT4-V is also precise in detecting micro-expressions. Its ability to process subtle and transient facial expressions enables it to accurately capture these fleeting emotional cues, which are often difficult for humans to perceive.
(3) GPT4-V’s performance in general facial expression recognition is not as accurate. This limitation may be due to the complexity and variety of facial expressions, resulting in the challenges in capturing and analyzing them. Nevertheless, when GPT4-V is used to process thought chains, its accuracy in facial expression recognition improves significantly. This improvement suggests that incorporating additional contextual information is of great importance to recognize facial expressions.
@@
(5) GPT4-V can also integrate with task-related agents to handle more complex tasks, such as detecting subtle facial changes and estimating heart rate with signal processing. By leveraging Python’s powerful libraries and tools, GPT4-V can effectively process and analyze intricate facial data to derive valuable insights, such as heart rate estimation, which can further enhance its applications in mental health monitoring and virtual human companion systems.
@@ 3.2 Tool call and processing
GPT-4V is one of the state-of-the-art multimodal language models that has achieved remarkable success in various natural language processing tasks. However, it is not directly applicable to some complex tasks, such as remote photoplethysmography (rPPG) [9, 18, 35, 44, 45, 46, 19, 26, 34, 2]. rPPG is a non-invasive technique used to measure heart rate and respiratory rate from facial videos. It has a wide range of applications in healthcare, entertainment, and human-computer interaction.
@@ 4 Further Discussion
GPT4 is a powerful language model that has shown remarkable success in various natural language processing tasks. However, it faces several challenges in other domains, such as facial expression recognition, emotion recognition, complex emotion recognition, non-contact physiological measurement, and authenticity detection.
Emotion recognition is the process of identifying and classifying emotions based on physiological signals, facial expressions, and speech patterns. GPT4 has shown promising results in this task; however, it requires a large amount of training data and may not generalize well to new datasets. To overcome this limitation, future research can focus on developing transfer learning techniques that enable GPT4 to learn from smaller datasets and generalize to new datasets.
Non-contact physiological measurement involves measuring physiological signals, such as heart rate, respiratory rate, and blood pressure, without direct contact with the body [9, 18, 19, 35, 34, 2]. GPT4 faces difficulty in this task due to its limited ability to process and interpret physiological signals accurately. To overcome this limitation, future research can focus on developing new technologies that can capture physiological signals accurately and integrate them with GPT4.
Deception detection involves identifying and verifying the authenticity of a person or an object. GPT4 faces difficulty in this task due to its limited ability to process and interpret visual and audio information accurately. To overcome this limitation, future research can explore ways to integrate GPT4 with computer vision and audio processing techniques to improve authenticity detection accuracy.
In conclusion, GPT4 faces several challenges in non-language tasks, such as facial expression recognition, emotion recognition, complex emotion recognition, non-contact physiological measurement, and deception detection. Future research can focus on developing new techniques to enhance GPT4’s ability to process and integrate multimodal data and improve its accuracy and efficiency in these tasks.
@@ 5 Conclusion
In this paper, we have discussed the challenges that GPT4 faces in non-language tasks, such as facial expression recognition, emotion recognition, complex emotion recognition, non-contact physiological measurement, and authenticity detection. While GPT4 has shown remarkable success in various natural language processing tasks, it faces limitations in these domains due to its limited ability to process and interpret visual and audio information accurately. To overcome these limitations, future research can focus on developing new techniques to enhance GPT4’s ability to process and integrate multimodal data and improve its accuracy and efficiency in these tasks. This may involve integrating GPT4 with computer vision and audio processing techniques, developing transfer learning techniques, exploring new sensor technologies, and improving the quality and quantity of training data. By addressing these challenges, GPT4 has the potential to significantly advance the fields of facial expression recognition, emotion recognition, complex emotion recognition, non-contact physiological measurement, and authenticity detection, and open up new avenues for research and application in these domains.

### 10 Apr 2024 v2

@@ Abstract
Multimodal large language models (MLLMs) are designed to process and integrate information from multiple sources, such as text, speech, images, and videos. Despite its success in language understanding, it is critical to evaluate the performance of downstream tasks for better human-centric applications. This paper assesses the application of MLLMs with 5 crucial abilities for affective computing, spanning from visual affective tasks and reasoning tasks. The results show that GPT-4V has high accuracy in facial action unit recognition and micro-expression detection while its general facial expression recognition performance is not accurate. We also highlight the challenges of achieving fine-grained micro-expression recognition and the potential for further study and demonstrate the versatility and potential of GPT-4V for handling advanced tasks in emotion recognition and related fields by integrating with task-related agents for more complex tasks, such as heart rate estimation through signal processing. In conclusion, this paper provides valuable insights into the potential applications and challenges of MLLMs in human-centric computing. Our interesting example is in https://github.com/EnVision-Research/GPT4Affectivity.
@@ 1 Introduction
The development of multimodal large language models (MLLMs) has been a topic of growing interest in recent years [23, 12, 24, 44, 22]. MLLMs are designed to process and integrate information from multiple modalities, such as text, speech, images, and videos. The development of these models has been driven by the need to improve the accuracy and efficiency of various tasks, such as affective computing, sentiment analysis, and natural language understanding.
MLLMs have shown great promise in improving the accuracy and robustness of affective computing systems [38, 28]. These models can process and integrate information from multiple modalities, such as facial expressions, speech patterns, and physiological signals, to infer emotional states accurately [12, 24, 44, 22] with significant implications for various applications, such as healthcare, education, and human-computer interaction.
[…]
Despite the rapid development of MLLMs, the need for standardized evaluation metrics is highlighted for accurate assessment. Different from the general language understanding evaluation benchmark that has been widely used to evaluate the performance of language models in NLP tasks [42, 41, 40], similar benchmarks to evaluate the performance of MLLMs for affective computing tasks are lacked, which is of great benefit to advance this field.
[…]
In this paper, we evaluate GPT-4V with 5 typical human-centric tasks, spanning from visual affective tasks and reasoning tasks. We summarize our findings as follows:
(1) GPT-4V is highly accurate in recognizing facial action units. This accuracy can be attributed to its advanced understanding of facial movements and their corresponding emotions, which allows it to effectively identify and analyze facial action units.
(2) GPT-4V is also precise in detecting micro-expressions. Its ability to process subtle and transient facial expressions enables it to accurately capture these fleeting emotional cues, which are often difficult for humans to perceive.
(3) GPT-4V’s performance in general facial expression recognition is not as accurate. This limitation may be due to the complexity and variety of facial expressions, resulting in the challenges in capturing and analyzing them. Nevertheless, when GPT-4V is used to process thought chains, its accuracy in facial expression recognition improves significantly. This improvement suggests that incorporating additional contextual information is of great importance to recognize facial expressions.
@@
(5) GPT-4V can also integrate with task-related agents to handle more complex tasks, such as detecting subtle facial changes and estimating heart rate with signal processing. By leveraging Python’s powerful libraries and tools, GPT-4V can effectively process and analyze intricate facial data to derive valuable insights, such as heart rate estimation, which can further enhance its applications in mental health monitoring and virtual human companion systems.
@@ 3.2 Tool call and processing
GPT-4V is one of the state-of-the-art multimodal language models that has achieved remarkable success in various natural language processing tasks. However, it is not directly applicable to some complex tasks, such as remote photoplethysmography (rPPG) [9, 19, 36, 45, 46, 47, 20, 27, 35, 2, 18]. rPPG is a non-invasive technique used to measure heart rate and respiratory rate from facial videos. It has a wide range of applications in healthcare, entertainment, and human-computer interaction.
@@ 4 Further Discussion
GPT-4V is a powerful language model that has shown remarkable success in various natural language processing tasks. However, it faces several challenges in other domains, such as facial expression recognition, emotion recognition, complex emotion recognition, non-contact physiological measurement, and authenticity detection.
Emotion recognition is the process of identifying and classifying emotions based on physiological signals, facial expressions, and speech patterns. GPT-4V has shown promising results in this task; however, it requires a large amount of training data and may not generalize well to new datasets. To overcome this limitation, future research can focus on developing transfer learning techniques that enable GPT-4V to learn from smaller datasets and generalize to new datasets.
Non-contact physiological measurement involves measuring physiological signals, such as heart rate, respiratory rate, and blood pressure, without direct contact with the body [9, 19, 20, 36, 35, 2]. GPT-4V faces difficulty in this task due to its limited ability to process and interpret physiological signals accurately. To overcome this limitation, future research can focus on developing new technologies that can capture physiological signals accurately and integrate them with GPT-4V.
Deception detection involves identifying and verifying the authenticity of a person or an object. GPT-4V faces difficulty in this task due to its limited ability to process and interpret visual and audio information accurately. To overcome this limitation, future research can explore ways to integrate GPT-4V with computer vision and audio processing techniques to improve authenticity detection accuracy.
In conclusion, GPT-4V faces several challenges in non-language tasks, such as facial expression recognition, emotion recognition, complex emotion recognition, non-contact physiological measurement, and deception detection. Future research can focus on developing new techniques to enhance GPT-4V’s ability to process and integrate multimodal data and improve its accuracy and efficiency in these tasks.
@@ 5 Conclusion
In this paper, we have discussed the challenges that GPT-4V faces in non-language tasks, such as facial expression recognition, emotion recognition, complex emotion recognition, non-contact physiological measurement, and authenticity detection. While GPT-4V has shown remarkable success in various natural language processing tasks, it faces limitations in these domains due to its limited ability to process and interpret visual and audio information accurately. To overcome these limitations, future research can focus on developing new techniques to enhance GPT-4V’s ability to process and integrate multimodal data and improve its accuracy and efficiency in these tasks. This may involve integrating GPT-4V with computer vision and audio processing techniques, developing transfer learning techniques, exploring new sensor technologies, and improving the quality and quantity of training data. By addressing these challenges, GPT-4V has the potential to significantly advance the fields of facial expression recognition, emotion recognition, complex emotion recognition, non-contact physiological measurement, and authenticity detection, and open up new avenues for research and application in these domains.